OpenStudy (18jonea):

You apply a force of 17 N to a wheel with a radius of 0.75 m. If the axle has a radius of 6 cm, what is the output force, assuming the machine operates under ideal conditions? 2.13 N 76.5 N 212.5 N 12.75 N @3mar

OpenStudy (3mar):

Do you have a graph or something?

OpenStudy (18jonea):

no i dont

OpenStudy (3mar):

so what you got?

OpenStudy (18jonea):

nothing but confused

OpenStudy (3mar):

really I need a figure because:: "a wheel with a radius of 0.75 m" and "the axle has a radius of 6 cm"

OpenStudy (18jonea):

i dont have a figure

OpenStudy (18jonea):

@3mar

OpenStudy (3mar):

@zepdrix

OpenStudy (3mar):

@Directrix

OpenStudy (3mar):

@mathmate

OpenStudy (18jonea):

this is the only thing i have in the notes

OpenStudy (mathmate):

@18jonea The figure/illustration means a lot to clarify the situation!

OpenStudy (18jonea):

so that helped?

OpenStudy (18jonea):

@mathmate

OpenStudy (3mar):

"The figure/illustration means a lot to clarify the situation!" THAT IS WHAT I WAS SAAAAAAAAAAAAAAAAYING IT MAKES SENSE NOWWWWWWW

OpenStudy (3mar):

The force at the wheel will be transefered as it is to the output axle. so the principle is \[F_1*r_1=F_2*r_\] where 1 stands for the moment at the wheel and 2 for the end axle 17*0.75=F_2*0.06 you find F_2

OpenStudy (3mar):

\[\huge F_1*r_1=F_2*r_2\]

OpenStudy (3mar):

Got it?

OpenStudy (18jonea):

12.75= f2 *.06

OpenStudy (18jonea):

12.75/.06= 212.5

OpenStudy (18jonea):

correct?

OpenStudy (18jonea):

@3mar

OpenStudy (3mar):

Correct! I had it 212.5 N also

OpenStudy (3mar):

Thank you for your patience.

OpenStudy (18jonea):

thank you!!!!

OpenStudy (3mar):

I really appreciate!

OpenStudy (3mar):

Thank you for the medal!

OpenStudy (3mar):

Any more questions?